# Edge AI: A Taxonomy, Systematic Review and Future Directions

### v1

@@ V-K Container Scaling
With the continuous development of cloud computing and container technology, container scaling has become an important means to ensure application performance, high availability, and resource optimization. This section will explore the strategies and practices of container scaling from two key perspectives: firstly, the scaling decisions of proactive and reactive, which exhibit different characteristics and advantages in dealing with load changes; next horizontal vertical and hybrid scaling strategies represent how effectively adjust container resources in different scenarios.
@@ V-K1 Proactive versus reactive scaling decisions
Proactive scaling decision: This method will use historical data of container load to train a specific AI model, through which future changes in container resource load can be perceived and predicted in advance [120]. It allows administrators or systems to automatically adjust resources to maintain optimal performance and efficiency. For example, this strategy can predict based on historical data that as long as it reaches 7pm or 8pm, the QPS of AI applications will significantly increase because everyone is off work, which is the entertainment time at night.
Reactive scaling decision: This is a strategy that utilizes third-party resource monitoring tools, such as Promethues [121], to make real-time decisions on the number of replicas and resource allocation in containers. The container orchestration tool determines whether to expand or dissolve based on the resource change data of the relevant containers in the monitoring tool. When the load increases, reactive scaling will start adding containers; When the load decreases, it will decrease the number of containers [122]. The decision-making of reactive scaling is based on real-time load data. When training the model, the utilization of GPU and GPU memory inside the container may reach 80%-90%. At this time, the system will immediately detect the high utilization rate and scale up the capacity promptly.
@@ V-K2 Horizontal, Vertical and Hybrid scaling
Horizontal scaling: It is a way to cope with load changes by increasing or decreasing the number of container instances [123, 124] (such as Pods, container groups, etc.). It can respond very quickly to load changes and adjust overall processing power by adding or removing container instances. Each container instance is independent and has good fault isolation, a fault in one instance will not affect other instances. Horizontal scaling is very suitable for scenarios with stateless services and the need to handle a large number of concurrent requests.
@@
Vertical scaling: It is adjusting the processing power of a single container instance by increasing or decreasing its resource allocation (such as CPU, memory, storage, etc.). It does not require managing multiple container instances, only adjusting the resource allocation of a single instance and can accurately adjust resource allocation based on actual load conditions, avoiding resource waste. Vertical scaling is suitable for stateful services[125]. However, this scaling strategy also has its drawbacks, as it poses a challenge to the computing and storage capabilities of individual machines.
@@
Hybrid scaling: Hybrid scaling is a scaling strategy that combines both horizontal scaling and vertical scaling [84]. Based on the load characteristics and requirements of the application, use both horizontal and vertical scaling methods to optimize resource allocation and performance. Being able to flexibly choose scaling methods based on different scenarios and needs, and combining horizontal and vertical methods can more effectively utilize resources, and improve application performance and stability.
@@ VI-A Infrastructure
In summary, cloud computing, fog computing, and edge computing have their own focus in Edge AI, and these three together build a complete edge intelligence ecosystem. Cloud computing provides powerful computing and storage resources, fog computing emphasizes real-time and low latency at the edge of the network, and edge computing enables real-time data processing and decision-making closest to the end device. These three work together to provide comprehensive support for the development of AI at the edge.
@@ VI-B Application
flexibility:Since all functional modules of the monolithic architecture run inside the same application, if we need to modify a module, we may have to recompile and deploy the entire application, so the monolithic architecture is not conducive to modularity and independent development, that is, it is less flexible; Each microservice in the corresponding microservices architecture can be deployed, scaled, and updated independently, making it easy to develop and maintain independently, thus increasing the flexibility of the overall application.
performance and resource utilization:Monolithic architectures may have resource contention and performance bottlenecks because all modules share the same process and resources, but from a resource utilization perspective, monolithic architectures may make more efficient use of resources because they do not require additional communication and management overhead; The microservices architecture, on the other hand, can independently deploy and scale the corresponding microservices according to the needs, thereby improving the performance of the application. From the perspective of resource utilization, services in a microservice fabric need to communicate with each other, which may increase the latency and bandwidth consumption of the system, and the microservice system may require more resources to manage and run due to the need to maintain multiple services.
deployment and scalability:Monolithic architectures are typically simple and easy to implement and deploy, but they often lack scalability to cope with frequently changing requirements; The corresponding microservices architecture, while more complex to deploy and often requires additional development and management efforts, scales flexibly and allows services to be added or removed quickly as needed.
@@ VI-C IoT Use Cases
There are significant differences between static and dynamic user use cases for IoT in edge AI. Static IoT user use cases typically involve the collection and analysis of static data, with relatively low data flow and frequency, low real-time requirements, and less variation in the location of the device. In contrast, dynamic IoT user use cases involve the collection and processing of real-time data, with higher data flow and frequency, higher real-time requirements, and the location of devices that can change frequently. In addition, static IoT use cases typically have relatively low energy consumption, while dynamic IoT use cases may consume more energy due to continuous data transmission and processing. Static IoT use cases tend to be more centralized when it comes to data processing and decision-making, while dynamic IoT use cases may require more distributed processing and decision-making. We give the specific differences between the two in Table IV:
TABLE IV: Comparison of IoT user use cases
Different use cases Static IoT user use cases Dynamic IoT user use cases
Data traffic Low High
Real-time Low High
Changes in location Low High
Energy consumption Low High
Data processing and decision-making Centralization Distributed
@@ VI-D Methods
For the four main AI methods, heuristics[127], meta-heuristics[74], machine learning[128], and deep reinforcement learning[129], we will compare them from the perspectives of applicable scenarios and problem complexity, data scale and training cost, real-time requirements and resource consumption, and generalization. Fig. 9 shows the specific comparison:
@@ VI-E Resource Management
With respect to the methods of resource provisioning, resource allocation, application placement, and workload distribution and prediction in edge AI resource management[130], we will further describe the relationship between these methods in detail, as shown in Fig. 10:
@@ VI-F ML Model Sizing
Regarding the deployment of AI models on edge devices, we generally have two deployment methods: compressed model [131, 132] and full model[133], and we compare these two methods from five aspects: model size, inference speed, accuracy, training and deployment cost, and application scenario, as shown in Table V:
TABLE V: Comparison of different deployment methods
Model Full model Compressed model
@@ VI-G Heterogeneity
The different types of heterogeneity[134] involved in edge AI deployment mainly include computing heterogeneity, hardware heterogeneity, and platform heterogeneity, as shown in Table VI:
TABLE VI: Comparison of different types of heterogeneity
@@ VI-H Security
Starting from the security of deploying edge devices, we mainly consider platform security, host security, and network security[135]. As shown in Fig. 11, we will compare the roles of the three in detail:
@@ VI-J Container Migration
There are four main types of container migration[136, 137] in edge AI: stateful migration and stateless migration, intra-cluster migration and inter-cluster migration, cloud/fog/edge migration, virtual migration and real-world testbed migration. We have made a detailed comparison of the different migration methods, as shown in Fig. 13:
@@ VI-K Container Scaling
Regarding container scaling[125] in edge AI, there are two main ways to actively scale and passively scale from the perspective of system response. From the expansion mode, there are mainly horizontal expansion, vertical expansion and hybrid expansion. As shown in Fig. 14, we make a detailed comparison of these two categories:
@@ VII Analysis and Result Outcomes
The survey enriches with various prospects associated with Edge empowered AI such as infrastructure support, IoT use cases, resource management strategies, security concerns and many more in the form of various state-of-the-art studies. The authors have systematically reviewed numerous articles in order to understand the prevailing the status of Edge AI in distinctive domains along with intelligent paradigms like ML and DL. There are lots of works going this direction to improve the life style of peoples and solve the real time problems. Hence, this section signifies the importance of our work referred in the form of year-wise papers, publication count, type of implementation (Simulation or experimental-based) and nevertheless QoS parameters addressed. Fig. 15 presents the year-wise analysis of related work carried in the form of number of papers referred from each year. The taxonomy of our study has been proposed with reference to articles from year 2015 to 2024. As depicted from Fig. 15, it is concluded that major chunk of the referred articles is recent and are from the year 2023. This clearly illustrates the fact that our survey includes the latest work done by the researchers.
@@
Apart from that, we rigorously reviewed the publication based statistics for the extensive study conducted highlighting its importance in real-time data processing. In total 1253 articles were collected during the data collection phase from various sources such as IEEE, ACM, Wiley, Science Direct, Taylor & Francis and Springer. Afterwards, filtering stage excluded collected articles based upon redundancy and inclusion and Exclusion criteria. The final stage comprises articles which the authors believe contributed the most towards shaping up the survey as depicted in Fig. 16 . Furthermore, the articles have been thoroughly reviewed and divided into 4 categories: review, Systematic Literature Review (SLR), implementation (simulation-based), implementation (Real/Testbed-based) and book. Fig. 17 illustrates that the major portion of the articles referred, are based on implementation (simulation-based), which signifies the fact that the implication of edge empowered-AI is yet to be tested on real-life IoT based use cases. Several real time work have been proposed by the researchers in recent years to improve the IoT applications based architecture using intelligent paradigm like ML, DL and reinforcement learning. This highlights a potential research direction for future studies to explore and validate edge-empowered AI in practical, real-world IoT environments. In addition, this article will motivate the researches to propose the novel solution to improve society 5.0.
@@ VIII Future Research Directions
Edge AI is continuously evolving and showing potential across various domains, offering numerous opportunities for innovation and improvement. This section examines critical future research directions that promise to enhance the capabilities and applications of Edge AI. These directions include optimizing energy use, strengthening security, and integrating with next-generation networks like 6G, highlighting the transformative impact of Edge AI across multiple sectors as shown in Fig. 18.
@@ VIII-A AI-driven Edge Data
AI-equipped edge devices can process data locally, providing real-time insights and actions without cloud transmission delays, crucial for applications like predictive maintenance in manufacturing. Using video, audio, and sensor data for anomaly detection, these devices enhance security in smart cities by identifying threats in real-time [138]. Additionally, AI-driven edge devices optimize energy usage by adjusting to weather, occupancy, and usage patterns in real-time [139]. Low latency and advanced privacy in AI model inference on the network edge benefit applications such as smart vehicles and drones [140]. The communication paradigm in Edge AI systems, designed to aggregate locally computed partial gradients at a central node, influences the performance of these edge nodes [140].
@@ VIII-B Energy
Optimizing energy use in AIoT systems through intelligent edge computing need to focus on sophisticated algorithms for dynamic energy distribution and task scheduling, ML for workload management, and designing low-power hardware [141]. Improvements in communication protocols and integrating renewable energy sources will enhance efficiency and scalability. Literature [142, 143] reported that AI techniques for managing renewable energy sources has been investigated, emphasizing advanced ML models for accurate forecasting and optimizing energy storage and grid integration. Implementing edge AI will reduce latency and enable real-time decision-making, increasing system responsiveness and resilience [144]. It has been identified that incorporating edge AI in the Internet of Energy (IoE), highlighting secure edge computing, blockchain for data security, lightweight AI algorithms, standardization for interoperability, and 5G networks for low-latency communication [75, 145, 146]. Advances in federated edge AI and DRL will optimize energy distribution, enhancing the efficiency and resilience of IoE systems [147].

### v2

@@ V-K Container Scaling
With the continuous development of cloud computing and container technology, container scaling has become an important means to ensure application performance, high availability, and resource optimization. This section will explore the strategies and practices of container scaling from two key perspectives: firstly, the scaling decisions of proactive and reactive, which exhibit different characteristics and advantages in dealing with load changes; next horizontal vertical and hybrid scaling strategies represent how effectively container resources are in different scenarios.
@@ V-K1 Proactive versus reactive scaling decisions
(i) Proactive scaling decision: This method will use historical data of container load to train a specific AI model, through which future changes in container resource load can be perceived and predicted in advance [124]. It allows administrators or systems to automatically adjust resources to maintain optimal performance and efficiency. For example, this strategy can predict based on historical data that as long as it reaches 7 pm or 8 pm, the QPS of AI applications will significantly increase because everyone is off work, which is the entertainment time at night.
(ii) Reactive scaling decision: This is a strategy that utilizes third-party resource monitoring tools, such as Promethues [125], to make real-time decisions on the number of replicas and resource allocation in containers. The container orchestration tool determines whether to expand or dissolve based on the resource change data of the relevant containers in the monitoring tool. When the load increases, reactive scaling will start adding containers; When the load decreases, it will decrease the number of containers [126]. The decision-making of reactive scaling is based on real-time load data. When training the model, the utilization of GPU and GPU memory inside the container may reach 80%-90%. At this time, the system will immediately detect the high utilization rate and scale up the capacity promptly.
@@ V-K2 Horizontal, Vertical and Hybrid scaling
(i) Horizontal scaling: It is a way to cope with load changes by increasing or decreasing the number of container instances [127, 128] (such as Pods, container groups, etc.). It can respond very quickly to load changes and adjust overall processing power by adding or removing container instances. Each container instance is independent and has good fault isolation, a fault in one instance will not affect other instances. Horizontal scaling is very suitable for scenarios with stateless services and the need to handle a large number of concurrent requests.
@@
(ii) Vertical scaling: It is adjusting the processing power of a single container instance by increasing or decreasing its resource allocation (such as CPU, memory, storage, etc.). It does not require managing multiple container instances, only adjusting the resource allocation of a single instance and can accurately adjust resource allocation based on actual load conditions, avoiding resource waste. Vertical scaling is suitable for stateful services[129]. However, this scaling strategy also has its drawbacks, as it poses a challenge to the computing and storage capabilities of individual machines.
@@
(iii) Hybrid scaling: Hybrid scaling is a scaling strategy that combines both horizontal scaling and vertical scaling [88]. Based on the load characteristics and requirements of the application, use both horizontal and vertical scaling methods to optimize resource allocation and performance. Being able to flexibly choose scaling methods based on different scenarios and needs, and combining horizontal and vertical methods can more effectively utilize resources, and improve application performance and stability.
@@ VI-A Infrastructure
In summary, cloud computing, fog computing, and edge computing have their own focus on Edge AI, and these three together build a complete edge intelligence ecosystem. Cloud computing provides powerful computing and storage resources, fog computing emphasizes real-time and low latency at the edge of the network, and edge computing enables real-time data processing and decision-making closest to the end device. These three work together to provide comprehensive support for the development of AI at the edge.
@@ VI-B Application
(i) Flexibility: Since all functional modules of the monolithic architecture run inside the same application, if we need to modify a module, we may have to recompile and deploy the entire application, so the monolithic architecture is not conducive to modularity and independent development, that is, it is less flexible; Each microservice in the corresponding microservices architecture can be deployed, scaled, and updated independently, making it easy to develop and maintain independently, thus increasing the flexibility of the overall application.
(ii) Performance and resource utilization: Monolithic architectures may have resource contention and performance bottlenecks because all modules share the same process and resources, but from a resource utilization perspective, monolithic architectures may make more efficient use of resources because they do not require additional communication and management overhead; The microservices architecture, on the other hand, can independently deploy and scale the corresponding microservices according to the needs, thereby improving the performance of the application. From the perspective of resource utilization, services in a microservice fabric need to communicate with each other, which may increase the latency and bandwidth consumption of the system, and the microservice system may require more resources to manage and run due to the need to maintain multiple services.
(iii) Deployment and scalability: Monolithic architectures are typically simple and easy to implement and deploy, but they often lack scalability to cope with frequently changing requirements; The corresponding microservices architecture, while more complex to deploy and often requires additional development and management efforts, scales flexibly and allows services to be added or removed quickly as needed.
@@ VI-C IoT Use Cases
IoT use cases can be divided into static and dynamic in terms of user mobility. As shown in Fig. 9 2, we will show the role of edge AI in two different IoT use cases from different perspectives.
Figure 9: Comparison of different IoT Use Cases
@@ VI-D Methods
For the four main AI methods, heuristics[131], meta-heuristics[78], machine learning[132], and deep reinforcement learning[133], we will compare them from the perspectives of applicable scenarios and problem complexity, data scale and training cost, real-time requirements and resource consumption, and generalization. Fig. 10 shows the specific comparison.
@@ VI-E Resource Management
With respect to the methods of resource provisioning, resource allocation, application placement, and workload distribution and prediction in edge AI resource management[134], we will further describe the relationship between these methods in detail, as shown in Fig. 11.
@@ VI-F ML Model Sizing
Regarding the deployment of AI models on edge devices, we generally have two deployment methods: reduced model [135, 136] and full model[137], and we compare these two methods from five aspects: model size, inference speed, accuracy, training and deployment cost, and application scenario, as shown in Table IV.
TABLE IV: Comparison of different deployment methods
Model Full model Reduced model
@@ VI-G Heterogeneity
The different types of heterogeneity[138] involved in edge AI deployment mainly include computing heterogeneity, hardware heterogeneity, and platform heterogeneity, as shown in Table V.
TABLE V: Comparison of different types of heterogeneity
@@ VI-H Security
Starting from the security of deploying edge devices, we mainly consider platform security, host security, and network security[139]. As shown in Fig. 12, we will compare the roles of the three in detail.
@@ VI-J Container Migration
There are four main types of container migration[140, 141] in edge AI: stateful migration and stateless migration, intra-cluster migration and inter-cluster migration, cloud/fog/edge migration, virtual migration and real-world testbed migration. We have made a detailed comparison of the different migration methods, as shown in Fig. 14.
@@ VI-K Container Scaling
Regarding container scaling[129] in edge AI, there are two main ways to actively scale and passively scale from the perspective of system response. From the expansion mode, there are mainly horizontal expansion, vertical expansion and hybrid expansion. As shown in Fig. 15, we make a detailed comparison of these two categories.
@@ VII Analysis and Result Outcomes
The survey enriches various prospects associated with Edge-empowered AI such as infrastructure support, IoT use cases, resource management strategies, security concerns and many more in the form of various state-of-the-art studies. The authors have systematically reviewed numerous articles in order to understand the prevailing status of Edge AI in distinctive domains along with intelligent paradigms like ML and DL. There are lots of works going in this direction to improve the lifestyle of people and solve real-time problems. Hence, this section signifies the importance of our work referred to in the form of year-wise papers, publication count, type of implementation (Simulation or experimental-based) and nevertheless QoS parameters addressed. Fig. 16 presents the year-wise analysis of related work carried out in the form of a number of papers referred from each year. The taxonomy of our study has been proposed with reference to articles from year 2015 to 2024. As depicted in Fig. 16, it is concluded that a major chunk of the referred articles is recent and are from the year 2023. This clearly illustrates the fact that our survey includes the latest work done by the researchers.
@@
Apart from that, we rigorously reviewed the publication-based statistics for the extensive study conducted highlighting its importance in real-time data processing. In total 1253 articles were collected during the data collection phase from various sources such as IEEE, ACM, Wiley, Science Direct, Taylor & Francis and Springer. Afterwards, the filtering stage excluded collected articles based on redundancy and inclusion and Exclusion criteria. The final stage comprises articles that the authors believe contributed the most towards shaping up the survey as depicted in Fig. 17. Furthermore, the articles have been thoroughly reviewed and divided into 4 categories: review, Systematic Literature Review (SLR), implementation (simulation-based), implementation (Real/Testbed-based) and book. Fig. 18 illustrates that the major portion of the articles referred, based on implementation (simulation-based), which signifies the fact that the implication of edge-empowered AI is yet to be tested on real-life IoT-based use cases. Several real-time works have been proposed by the researchers in recent years to improve the IoT applications-based architecture using intelligent paradigms like ML, DL and reinforcement learning. This highlights a potential research direction for future studies to explore and validate edge-empowered AI in practical, real-world IoT environments. In addition, this article will motivate the researchers to propose novel solutions to improve society 5.0.
@@ VIII Future Research Directions
Edge AI is continuously evolving and showing potential across various domains, offering numerous opportunities for innovation and improvement. This section examines critical future research directions that promise to enhance the capabilities and applications of Edge AI as summarized in Figure 19. These directions include optimizing energy use, strengthening security, and integrating with next-generation networks like 6G, highlighting the transformative impact of Edge AI across multiple sectors.
@@ VIII-A AI-driven Edge Data
AI-equipped edge devices process data locally, enabling real-time decision-making without the latency caused by cloud transmission. This is essential for latency-sensitive applications such as predictive maintenance, real-time video analytics, and autonomous systems. In predictive maintenance, for instance, AI models deployed at the edge process sensor data from machines to predict failures and schedule repairs before breakdowns occur. These systems often rely on time-series forecasting models, such as Long Short-Term Memory (LSTM) networks, which are optimized for real-time edge inference through techniques like pruning and quantization [142].
Using video, audio, and sensor data for anomaly detection, edge AI enhances security in smart cities by identifying threats in real-time [143]. Anomaly detection often employs CNNs or RNNs to detect irregularities in sensor streams. CNNs, such as MobileNet, use depthwise separable convolutions to reduce the number of operations while preserving spatial hierarchies, making them suitable for resource-constrained edge devices. To further optimize performance, techniques like model compression and pruning are used to reduce the model size without significantly compromising accuracy [142]. RNN-based architectures, including Gated Recurrent Units (GRUs), are also deployed at the edge for detecting anomalies in sequential data, such as traffic patterns or environmental sensors.
Reinforcement learning algorithms are widely used to manage energy consumption at the edge, particularly for dynamic resource allocation. In HVAC systems, RL models learn optimal policies to balance energy consumption and occupant comfort by interacting with the environment and receiving feedback in the form of rewards [144]. In these systems, deep Q-learning (DQN) is used to handle large state-action spaces efficiently. Edge devices also deploy on-policy methods like Proximal Policy Optimization (PPO), which allow continuous adjustments based on real-time data. These models are further optimized through techniques like experience replay, which reduces memory usage and computation load, critical for edge deployments.
Edge AI plays a pivotal role in smart vehicles and drones, where real-time sensor data processing is crucial for navigation and decision-making [145]. Autonomous vehicles rely on edge AI to process data from LIDAR, RADAR, and cameras using sensor fusion techniques, which combine multiple sensor inputs to improve accuracy and robustness [146]. For instance, Kalman filtering is employed to integrate noisy sensor measurements, while CNNs perform object detection and classification. Edge AI uses these techniques to make split-second decisions, such as obstacle avoidance, without cloud dependencies. Privacy-preserving models like federated learning are also crucial in this context, enabling local data processing on vehicles while sharing only model updates, ensuring that sensitive location data remains private. Advanced techniques such as differential privacy and homomorphic encryption are integrated into FL to protect against data leakage during model updates.
However, deploying AI at the edge presents significant challenges, particularly in terms of computational limitations and energy constraints. Techniques like model quantization, where neural network weights are reduced from 32-bit floating point to 8-bit integers, help decrease the model size and improve inference speed [142]. Moreover, hardware-specific optimizations, such as leveraging the parallelism of Tensor Processing Units (TPUs) or Graphics Processing Units (GPUs), play a crucial role. For example, Google’s Edge TPU accelerates inferencing tasks with high energy efficiency, while Nvidia’s Jetson platform provides scalable computing power for more complex tasks.
In edge computing, the communication paradigm plays a critical role in system performance, especially in distributed learning models like Federated Learning [145]. FL enables devices to compute local updates on their own datasets and only transmit model gradients, thereby reducing communication overhead. Challenges such as non-IID (non-independent, identically distributed) data among edge nodes, which can lead to model biases, are addressed through methods like FedProx, which adds a regularization term to prevent drastic divergence from the global model. Additionally, communication-efficient techniques such as gradient sparsification, where only significant gradients are transmitted, and asynchronous updates ensure that edge nodes can update models independently without waiting for synchronization, thus improving efficiency in bandwidth-constrained environments.
The convergence of edge AI and advanced hardware solutions continues to drive innovations in AI deployments. Custom AI accelerators, such as Google’s Edge TPU, are optimized for low-power inference tasks, supporting applications that require high-speed processing, such as object detection in real-time video feeds [146]. NVIDIA’s Jetson platform, on the other hand, provides a scalable solution for more compute-intensive tasks, such as deep learning-based robotics or autonomous navigation. These platforms support parallelized inference operations, maximizing throughput while minimizing energy consumption, making them ideal for edge environments. Future research directions include developing neuromorphic hardware, which mimics biological neural networks, offering significant reductions in energy consumption while maintaining high processing speeds.
@@ VIII-B Energy
Optimizing energy use in AIoT systems through intelligent edge computing requires focusing on sophisticated algorithms for dynamic energy distribution, task scheduling, ML for workload management, and the design of low-power hardware [147]. For example, dynamic energy distribution can be handled using RL-based approaches where the system continuously learns and adapts to energy usage patterns. Techniques such as DQN and Multi-agent Reinforcement Learning (MARL) are effective for decentralized energy optimization, allowing edge devices to work autonomously while coordinating with other nodes in the network to maximize efficiency. MARL models allow each device to act as an independent agent, optimizing energy at both local and global levels by sharing information across the network [148]. Furthermore, task scheduling in AIoT systems can benefit from heuristic algorithms like Genetic Algorithms (GA) and Particle Swarm Optimization (PSO), which provide near-optimal solutions for task allocation in energy-constrained environments. These methods are computationally efficient, making them well-suited for edge devices with limited resources [149, 150].
Improvements in communication protocols and integration of renewable energy sources will further enhance system efficiency and scalability. Protocols such as Low-Power Wide Area Networks (LPWANs), including LoRaWAN and NB-IoT, are particularly useful for AIoT systems as they allow for long-range communication with minimal power consumption. These protocols, when combined with AI models running at the edge, enable real-time data exchange between devices while conserving energy. Additionally, optimizing the scheduling of data transmissions based on energy availability or demand-response signals can significantly reduce communication overhead in energy-constrained environments [151].
Literature [152, 153] reported that AI techniques for managing renewable energy sources have been investigated, emphasizing advanced ML models for accurate forecasting and optimizing energy storage and grid integration. For instance, LSTM networks and Autoregressive Integrated Moving Average (ARIMA) models have been widely applied in energy forecasting tasks. LSTMs are especially effective in capturing the temporal dependencies in renewable energy generation data (e.g., solar and wind), allowing for more accurate predictions of energy availability. These forecasts can then be used to optimize the allocation of tasks across the grid, improving the balance between energy demand and supply. In energy storage, AI models like Gradient Boosting Machines (GBM) and Support Vector Machines (SVM) have been applied to optimize the charge-discharge cycles of batteries, maximizing the longevity and efficiency of storage systems [154]. Integrating these AI models with grid control systems enables real-time decision-making to manage fluctuations in energy generation and consumption effectively.
Implementing edge AI will reduce latency and enable real-time decision-making, increasing system responsiveness and resilience [155]. In edge computing environments, AI models can be used to make decisions locally without needing to transmit large amounts of data to a central cloud, which reduces the overall latency and bandwidth usage. This is especially beneficial in scenarios where immediate action is required, such as energy demand-response events or real-time fault detection in smart grids. Edge AI, combined with lightweight models such as MobileNet or TinyML, can process sensor data in real time, detecting anomalies or optimizing energy usage without draining significant computational resources. This local processing not only reduces decision latency but also enhances the system’s resilience to network disruptions.
It has been identified that incorporating edge AI in the Internet of Energy (IoE) presents unique opportunities, particularly in areas like secure edge computing, blockchain for data security, lightweight AI algorithms, standardization for interoperability, and 5G networks for low-latency communication [79, 156, 32]. Blockchain technology offers decentralized solutions for secure energy trading and data management within IoE networks. The integration of blockchain with edge AI can ensure the transparency and immutability of energy transaction data while minimizing the energy overhead typically associated with blockchain’s consensus mechanisms. By using energy-efficient consensus protocols like Proof of Stake (PoS) or Proof of Authority (PoA), IoE systems can maintain security without incurring significant computational costs. Additionally, 5G and the upcoming 6G networks offer ultra-low latency and high-speed connectivity, enabling real-time energy optimization by allowing AI models to dynamically offload tasks between the edge and cloud based on energy constraints. Network slicing within 5G/6G networks ensures that energy-critical tasks receive prioritized resources, improving the overall performance and responsiveness of the IoE system.
Advances in federated edge AI and DRL will optimize energy distribution, enhancing the efficiency and resilience of IoE systems [157]. Federated Learning allows edge devices to collaboratively train a global model without exchanging raw data, thereby preserving privacy and reducing communication overhead. In energy management, FL can be adapted to enable energy-aware federated updates, where devices with limited energy resources reduce their participation in the model training process[154]. This approach ensures that the overall system remains efficient, even when certain devices operate under low-power conditions. DRL further improves energy optimization by learning policies for real-time control of energy resources, such as dynamically adjusting the operation of energy storage systems or scheduling energy-consuming tasks during periods of high renewable energy generation. These algorithms, when combined with edge AI, enable IoE systems to respond to changes in energy supply and demand with minimal delay.